Write each of the following as decimals:
3 / 2 ?????????

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

3 / 2 = (2 + 1) / 2
=2 / 2 + 1 / 2
= 1 + 0.5
= 1.5

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve solve this question as follows :
 
3/2 = (1+2)/2 = 1/2 + 2/2
      = 0.5 + 1 = 1.5
 
Hope it helps!
Thank you
 
Hope it helps!
Thank you